OpenStudy (anonymous):

PLease help me figure this out step by step! thank u so much! (-10x^3+30x-20)/(-5x+5) equals?

OpenStudy (anonymous):

saw this exact same problem yesterday first factor out a \(-5\) top and bottom to get \[\frac{-5(2x^3-6x+4)}{-5(x-1)}=\frac{2x^3-6x+4}{x-1}\]

OpenStudy (anonymous):

then factor the numerator as \[(x-1)(x-1)(2x+4)\] and cancel to get \[(x-1)(2x+4)\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

nvrm got it